Flexible Overhead Budget

Leno Manufacturing Company prepared the following factory overhead cost budget for the Press Department for October of the current year, during which it expected to require 12,000 hours of productive capacity in the department:

Variable overhead cost:
   Indirect factory labor $88,800
   Power and light 5,040
   Indirect materials 27,600
      Total variable overhead cost $121,440
Fixed overhead cost:
   Supervisory salaries $42,500
   Depreciation of plant and equipment 26,720
   Insurance and property taxes 17,000
      Total fixed overhead cost 86,220
Total factory overhead cost $207,660

Assuming that the estimated costs for November are the same as for October, prepare a flexible factory overhead cost budget for the Press Department for November for 10,000, 12,000, and 14,000 hours of production. Round your interim computations to the nearest cent, if required. Enter all amounts as positive numbers.

Leno Manufacturing Company
Factory Overhead Cost Budget-Press Department
For the Month Ended November 30
Direct labor hours 10,000 12,000 14,000
Variable overhead cost:
Indirect factory labor $ $ $
Power and light
Indirect materials
Total variable factory overhead $ $ $
Fixed factory overhead cost:
Supervisory salaries $ $ $
Depreciation of plant and equipment
Insurance and property taxes
Total fixed factory overhead $ $ $
Total factory overhead cost $ $ $
Answer #1
Leno Manufacturing Company
Factory Overhead Cost Budget-Press Department
For the Month Ended November 30
Direct labor hours 10000 12000 14000
Variable overhead cost:
Indirect factory labor 74000 88800 103600
Power and light 4200 5040 5880
Indirect materials 23000 27600 32200
Total variable factory overhead 101200 121440 141680
Fixed factory overhead cost:
Supervisory salaries 42500 42500 42500
Depreciation of plant and equipment 26720 26720 26720
Insurance and property taxes 17000 17000 17000
Total fixed factory overhead 86220 86220 86220
Total factory overhead cost 187420 207660 227900
Workings:
Direct labor hours 10000 14000
Variable overhead cost:
Indirect factory labor =88800/12000*10000 =88800/12000*14000
Power and light =5040/12000*10000 =5040/12000*14000
Indirect materials =27600/12000*10000 =27600/12000*14000